OpenStudy (anonymous):

If A + B = 90°, sin A = p/q, cos B = r/s, the correct relation between p, q, r, s will be

OpenStudy (anonymous):

same compendendo and dividendo rule ahh

OpenStudy (anonymous):

sin^2+cos^2=1

ganeshie8 (ganeshie8):

is it given that ABC is a triangle or something ?

OpenStudy (anonymous):

nope:)i was blindly saying

ganeshie8 (ganeshie8):

A + B = 90 A = 90 - B sin(A) = sin(90-B) = cosB

ganeshie8 (ganeshie8):

sinA = cosB so simply equate : p/q = r/s